### v1

# Scaling Up On-Device LLMs via Active-Weight Swapping Between DRAM and Flash

Abstract.
Large language models (LLMs) are increasingly being deployed on mobile devices, but the limited DRAM capacity constrains the deployable model size. This paper introduces ActiveFlow, the first LLM inference framework that can achieve adaptive DRAM usage for modern LLMs (not ReLU-based), enabling the scaling up of deployable model sizes. The framework is based on the novel concept of active weight DRAM-flash swapping and incorporates three novel techniques: (1) Cross-layer active weights preloading. It uses the activations from the current layer to predict the active weights of several subsequent layers, enabling computation and data loading to overlap, as well as facilitating large I/O transfers. (2) Sparsity-aware self-distillation. It adjusts the active weights to align with the dense-model output distribution, compensating for approximations introduced by contextual sparsity. (3) Active weight DRAM-flash swapping pipeline. It orchestrates the DRAM space allocation among the hot weight cache, preloaded active weights, and computation-involved weights based on available memory. Results show ActiveFlow achieves the performance-cost Pareto frontier compared to existing efficiency optimization methods.
1. Introduction
Large language models (LLMs) are increasingly deployed on mobile and PC devices as integral system components, such as the on-device 3B Apple foundation model for Apple iOS (Apple, [n. d.]), the 3.82B Phi Silica for Windows (Microsoft, [n. d.]), and 3.35B Gemini Nano for Google’s Android (Team, 2024).
However, further scaling up the on-device LLM size is very difficult, with a key constraint of DRAM size. Due to power and area constraints, the DRAM size on mobile devices remains limited and difficult to increase, even across device upgrades (e.g., both iPhone 15 and iPhone 16 feature 8 GB DRAM). Furthermore, the available DRAM capacity is also determined by the co-active apps and OS processes remaining in DRAM simultaneously. Mobile OS can terminate an app under low available DRAM unless the app can reduce the memory usage(Xcode, 2024).
Goal. To enable the deployment of larger LLMs, it is essential to realize adaptive DRAM usage for LLM inference. That is, the inference process dynamically adapts to different available DRAM sizes while maintaining comparable model quality and inference speed. Mirroring the OS employs virtual memory to abstract physical limitations, this work aims for adaptive DRAM usage that is transparent to the user, creating the illusion that the entire model resides in DRAM.
Adaptive DRAM usage has been previously investigated for traditional non-autoregressive DNNs (e.g., CNN and Bert) through DRAM-Flash swapping (Li et al., 2024; Yi et al., 2023). However, the fundamental difference in workload characteristics hinders the direct application of these methods to LLMs. Existing techniques rely on the computation-intensive feature of traditional DNNs, so the current operator computation can overlap the loading of the next operator. While this overlap is present in the LLM prefilling stage, the significantly more time-consuming autoregressive decoding phase is bottlenecked by memory access. Consequently, realizing user-oblivious adaptive memory management for LLM inference necessitates minimizing Flash data loading to mitigate the substantial disparity between memory and Flash bandwidth (5$\sim$ 5$\times$ 5 × on mobile phones).
Fortunately, a unique characteristic of LLMs is contextual sparsity, where although the model itself is large, only a small subset of weights is actively used per token generation (Liu et al., 2023), which we term as active weights. Our upper-bound analysis (Fig. 2) shows that during each inference iteration, only ¡15% weights need to be activated to generate the same token.
Figure 1. The perplexity of LLaMA-3-8B model on various memory cost. Ours shows the best performance compared with SOTA quantization (Q), pruning (P) and contextual sparsity (SP). Each point on the scaling line means a sparsity ratio.
Challenges. This contextual sparsity inspires us explore the new opportunity of active weights swapping for adaptive memory usage. Unlike traditional per-operator swapping, active weight swapping introduces greater challenges: (1) How to accurately identify the active weights, given contextual sparsity is highly dynamic, varying cross tokens, layers and blocks. Misidentification could degrade model accuracy. (2) How to predict the active weights as early as possible, allowing for overlapping computation with loading, as well as efficient large I/O transfers, both of which are critical for performance.
Several works have explored contextual sparsity(Liu et al., 2023; Xue et al., 2024; Song et al., 2024b; Wang et al., 2024b; L et al., 2024; Mirzadeh et al., 2023; Song et al., 2024c, a), but gaps remain in addressing the challenges above. Some methods like Deja Vu (Liu et al., 2023), PowerInfer (Song et al., 2024b) and LLM in a flash (Alizadeh et al., 2023) use available ReLU-based models to generate zero activations and introduce additional predictors (GB memory cost) to forecast these zeros. However, modern LLMs used in productions (e.g., Llama) rarely use ReLU-based architecture due to its inferior accuracy (Touvron et al., 2023) (see Fig.14(b)). There are also works performing continued pre-training to transform available models to ReLU or ReLU-variant based, such as PowerInfer-2 (Xue et al., 2024), TurboSparse (Song et al., 2024c), ProSparse (Song et al., 2024a), and Q-Sparse (Wang et al., 2024b). These works require training on hundreds of billions of tokens and consume substantial hardware resources. TEAL (Liu et al., 2025a) proposes a training-free, magnitude-based sparsity method (see Fig. 3), where only activations above a threshold are computed. However, the active weights cannot be predicted, but only be identified after the input activation is ready. Additionally, the method is empirical, and there is no mechanism to compensate for the accuracy loss due to the potential misidentification of active weights. Therefore, current techniques fall short of achieving adaptive memory usage for LLMs.
Figure 2. The upper bound sparsity of LLaMa-2-70B model during decoding.
Figure 3. The ReLU-based sparsity and Top-K activation sparsity. We base our system on Top-K sparsity due to its broader applicability and higher accuracy.
Our work. This paper proposes ActiveFlow LLM inference framework. It can realize user-oblivious adaptive DRAM usage, in order to scale up the LLM sizes that can be deployed on mobile devices. Similar to TEAL, this paper utilizes magnitude-based, model-architecture-independent activation sparsity, to ensure the framework’s applicability to modern LLMs. Beyond that, ActiveFlow incorporates three novel techniques.
Firstly, Cross-layer active weight preloading. To address the sequential dependency issue of active weights with its input activation in order to enable computation and loading overlapping, we propose cross-layer active weight preloading. It creatively utilizes the current layer’s activation to pre-identify the next n layers’ active weights. It is based on the key obeservation that due to the widely used residual connection, the activation magnitude distribution across layers share significant similarity (¿80% shown in Fig. 4(a)). For the active weights that missed by pre-loading, ActiveFlow loads on-demand when the actual activation is ready.
Secondly, Sparsity-aware self-distillation. Even the magnitude-based activation sparsity empirically has shown the superior quality compared to other sparse methods (Liu et al., 2025b), it still introduces an approximation compared to the dense model. To compensate for the approximation, we propose sparse-aware self-distillation to adjust the active weights towards the dense-model output. The distillation improves both the sparsity ratio and model accuracy. The technique is inspired by and integrated with the quantization-aware self distillation (Du et al., 2024). Similar to this work, the self-distillation only needs several A100 GPU hours to train. The two methods can be used collaboratively for LLM deployment.
Thirdly, DRAM-flash active weight swapping pipeline. The pipeline reorganizes the data layout for the cross-layer preloading, and overlaps the active weight loading with the current layer computing. It also integrates a contextual hot active weight caching policy beyond naive swapping. The pipeline orchestrates the space allocations among the cache, preloaded active weights, and computation involved weights according to available memory.
We implement ActiveFlow and evaluate it on different mobile phones (OnePlus 12, Pixel 6, and Infinix Zero). Results (Fig. 1, more in Sec. 7) show that ActiveFlow achieves the inference performance-cost Pareto frontier among existing efficiency optimization methods, including state-of-the-art quantization (DB-LLM (Chen et al., 2024) and PB-LLM (Shang et al., 2023)), pruning (CPSP (Wang et al., 2024d) and RIA (Yi and Richtárik, 2025)), and contextual sparsity (Teal (Liu et al., 2025b)), demonstrating its practical value. Particularly, under the same model quality and speed, ActiveFlow reduces the DRAM usage by up to 40% for Llama 7B compared to llama.cpp. Under the same sparsity ratio, ActiveFlow can reduce memory by 2$\times$ compared to Teal. ActiveFlow is the first to successfully deploy the original Mixtural-8x7B 4bit model (Jiang et al., 2024) (no ReLU introduced) on a mid-range pixel-6 phone, achieveing 1.8 tokens/s with 2.9 GB memory cost.
To summarize, the contributions of this paper are:
  • We propose ActiveFlow, the first LLM inference system to enable user-oblivious adaptive DRAM usage through active weight swapping for modern general LLMs without ReLU dependency.
  • We propose the cross-layer active weights preloading to allow computation/loading overlapping and large I/O transfer.
  • We propose sparsity-aware self distillation to compensate the approximation introduced by sparsity.
  • We implement the end-to-end ActiveFlow. Results show it achieves the inference quality-cost Pareto frontier among existing optimization methods.
2. Motivation and Background
2.1. Upper Bound Analysis of Contextual Sparsity in LLMs
A specific feature of LLMs is contextual sparsity (Liu et al., 2023; Xue et al., 2024; Song et al., 2024b; Wang et al., 2024b; L et al., 2024; Mirzadeh et al., 2023; Song et al., 2024c), which means a small, context-dependent subset of total weights, that can generate the same output as the full model. We term this small subset of weights as active weight. Compared to the static sparsity from model pruning (Frantar and Alistarh, 2023; Sun et al., 2024), contextual sparsity dynamically selects different active weights for computation during each token generation, preserving the model’s overall capacity and adaptability. Contextual sparsity has also been empirically demonstrated to be compatible with model quantization (Wang et al., 2024b).
Since our techniques will be based on contextual sparsity, we first analyze the upper bound of this sparsity. We use a Llama-2-70B model to evaluate the amount of active weights required to generate the same token with full weights during the decoding process. The evaluation is conducted by incrementally removing unimportant weights for each decoded token by 1%. The important scores of weights are calculated by $S_{ij}=|W_{ij}|\cdot|X_{j}|$ , where $W_{ij}$ is an element of weight matrix and $X_{j}$ is an element of the input activation vector. As shown in Fig. 2, the results indicate that most tokens require less than 5% of the weights, with the maximum active weight being only 15%. This high level of sparsity shows a great potential for reduced inference cost.
Although the above results are promising, it is challenging to identify the active weights during inference, unless the weights are loaded and computed with activations. Consequently, some works (Liu et al., 2023; Song et al., 2024a) rely on ReLU-generated sparsity and propose extra predictors to estimate the sparsity, as illustrated in Fig. 3(a). These predictors are trained with calibration datasets, loaded into memory, and executed before performing per-layer LLM computations. However, the deployment cost of predictors is significant because (1) the datasets may not be suitable for real user data, (2) predictors require additional memory (at the GB level), and (3) they introduce extra computational overhead.
More recent works (Liu et al., 2025b; Wang et al., 2024b) propose magnitude-based activation sparsity, as shown in Fig. 3(b). We term this sparsity as Top-K sparsity following (Wang et al., 2024b). Only the activation elements with a magnitude above a threshold will be computed for each operator. Top-K sparsity demonstrates obvious advantages: 1) compatibility to modern non-ReLU LLMs; 2) applicability to all linear transformation operators rather than just FFN blocks; 3) no extra predictors needed.
These advantages motivate us to identify active weights for swapping based on Top-K activation sparsity.
2.2. Observation: Similarities in Cross-Layer Activations
Figure 4. The cross-layer input activation similarity of a LLaMA-2-7B model. (a) The attention input cosine similarity and top-k precision. (b) The value of activation before/after LayerNorm layer and average weights.
Figure 5. The simplified transformer layer structure of an LLM model. Residual connections pass the input of a block directly to output.
A key observation of this paper is that the input activations of the attention and MLP blocks in LLMs exhibit high cross-layer similarity. Fig. 4(a) uses the input activation of the attention block as an example to show the cosine similarity and top-k sparsity precision in each consecutive two layers in a Llama-2-7B model. Starting from the 3rd layer, the attention Q, K, V, and FFN gate and up operators exhibit over 95% similarity. Consequently, the top-k sparsity precision for these operators exceeds 80% cross layers.
The similarity is primarily due to the significant contribution of the residuals to the input activations. Fig.  5 shows a simplified transformer layer structure. The input activations are composed of the sum of two elements: the output activation of the previous block $F(X)$ and the residual $X$. The cross-layer similarity is because the residual values $X$ are larger than the output activation values $F(X)$. This difference in values arises from (1) the LayerNorm layer in the attention and MLP blocks, and (2) the weights magnitudes. As shown in Fig. 4(b), the LayerNorm reduces the activation magnitude by 50%. Additionally, the weight magnitude is smaller than the activation magnitude, resulting in a smaller calculation output.
The cross-layer input similarity motivates us for cross-layer preloading, which uses current layer’s activation to identify following layers’ active weights.
2.3. Observation: Contextual Hot Active Weights During Decoding
Figure 6. The selection probability of active weights in attention Q/K/V operators of Llama-2-7B model (under 50% contextual sparsity). Context level shows higher selection probability than task level. We only show the active weight with probability ¿ 0.7.
This section investigates the presence of hot active weights, i.e., the weights that are frequently selected across inference iterations during decoding. This investigation aims to identify opportunities for caching and more intelligent swapping strategies. Our observation is that contextual active weights exhibit high temporal locality across inference iterations during decoding, suggesting that caching hot active weights for higher cache hit rates.
As shown in Fig. 6, we conducted two levels of active weight selection frequency analysis: task level and context level. The task level counts the frequency with which weight channels are selected during the decoding process for all input contexts across a dataset (WikiText-2). In contrast, the context level counts the frequency of weight selection specifically for the decoding process of a given input context. Results show that hot weight selection probabilities on the context level exceeds 0.7, while the task level exceeds 0.5. The difference demonstrates the potentially improved cache hit and reduced loading cost by implementing a contextual cache management policy.
3. Cross-layer Active Weight Preloading
Figure 7. The flash read throughput of various IO chunk sizes on three devices with difference UFS capabilities.
To realize adaptive DRAM usage, two critical challenges for performance is: (1) whether the weight loading and computation can be overlapped to hide the flash loading overhead; (2) whether the I/O transfer can fully utilize the flash bandwidth. As shown in Fig. 7, the flash read throughput varies greatly with the chunk size of each I/O transfer. To achieve the peak flash throughput, the chunk size has to ¿64 KB. However, active weight from Top-K activation sparsity is in channel granularity, e.g., 4KB (see Fig. 3), and naive loading of the each active weight channel from flash can reduce the throughput from GB/s to MB/s.
However, current works including PowerInfer (Xue et al., 2024; Song et al., 2024b), LLM in Flash (Alizadeh et al., 2023) and Ripple (Wang et al., 2024a) only partially alleviated the problem. To enlarge the chunk size, they cluster co-active weight channels within the same block, and overlap each cluster loading and computation.
Our technique. To overcome the challenges, based on our key observation that cross-layer activations exhibit significant similarity, we propose the cross-layer active weight preloading. As shown in Fig. 8, while the computing of current layer, the next N layers’ active weights will be preloaded to DRAM simultaneously. We term these N layers as a layer group for preloading. The N is set based on the available DRAM, and the computing latency (N=4 can fully overlap the loading and computing in our evaluation). The preloading will include the active weights from all the operators in both Attention and FFN blocks. Different activations correspond to different parts of the weights being loaded. For example, Q, K, and V activations are only used to load $W_{q}$, $W_{k}$, and $W_{v}$, respectively.
Figure 8. Cross-layer active weight pre-loading. While the computing of current layer, the active weights of all the operators in the next N layers (layer group) will be preloaded based on the current activation. The missed active weights during preloading will be on-demand loaded after its actual activation is ready.
Figure 9. The reordered weights in a 4-layers group. The weight layout now is in the order of weight channel, layer, and operator type. By multi-layer weight reordering, the minimal loading chunk size is increased to improve the loading efficiency.
Since cross-layer activation similarity is not 100%, pre-loading can only load a portion of the necessary weights in advance. Any remaining weights that were not correctly pre-loaded are fetched through on-demand loading. This only takes $\sim 5\%$ of the total active weights.
[…]
Figure 10. The computing-loading overlap pipeline for LLM inference in an attention block after warming up.
Figure 11. The weight layout and flow of ActiveFlow.
4. Active Weight Swapping Pipeline
Building on the proposed cross-layer-group LLM weight loading and reordering techniques, we design a LLM computing-loading overlapping execution pipeline as shown in Fig. 10. The pipeline consists of four main operations: (1) Computing (C) – Performs the required computations. (2) Top-k (T) – Extracts the top-k mask from activations to determine the indices of the activated weight channels. (3) On-demand loading (L) – Loads weights for the current layer group. (4) Preloading (PL) – Preloads weights for the next layer group.
Fig. 11 demonstrates the weight layout and flow with the pipeline. The whole model resides in the flash with the cross-layer group layout. The current active weights, as well as the pre-loaded and cached weights store in the DRAM. The computation and loading are concurrently executed.
The overlapped LLM execution pipeline follows two key principles: (1) Maximize the overlap between loading and computing to minimize idle time (bubbles) to fully utilize the memory bandwidth and computing power simultaneously. (2) Maximize the cache hit rate on the sequence level. The challenge is how to accurately estimate the impact of system parameters, such as sparsity, memory cost and cache size on the accuracy and latency of model inference.
Table 1. The symbols of our system cost model.
Symbols Description
$sp$ sparsity of LLM
$hr$ average hit rate of weight cache
$si$ average similarity of cross-layer group
$BW_{mem}$ bandwidth of memory
$BW^{small}_{flash}$ bandwidth of small chunk reading from flash
$BW^{large}_{flash}$ bandwidth of large chunk reading from flash
$S_{m}$ Size of LLM
$S_{l}$ Size of a LLM layer
$N$ Layer number of a cross-layer group
$M$ Memory cost of pipeline
$M_{max}$ Memory budget
$M_{cl}$ Memory of a cross-layer group
$M_{cache}$ Memory of weight cache
$M_{kv}$ Memory of KV cache
$T_{decode}$ Decoding time of a token
$T_{load}$ Loading time of a cross-layer group
$T_{comp}$ Computing time of a cross-layer group
$T_{overlap}$ Overleaping time of two cross-layer groups
$T_{onload}$ On-demand loading time of a cross-layer groups
$T_{preload}$ Preloading time of a cross-layer groups
4.1. Elastic and Optimized LLM Execution
The goal of this technique is to determine the optimal system parameters, including LLM sparsity, layer number of a cross-layer group, and cache size, for a given mobile device (i.e., with specific computational power and memory budget) and a given LLM. The objective is to minimize system latency while respecting the memory constraint.
There is tradeoff between LLM sparsity, layer number of a cross-layer group and cache size on the inference metrics in terms of both latency and accuracy. Optimizing one metric could worsen another. To capture this, we define the following problem, with the memory cost as a hard constraint and the objective to minimize the decode latency. The related symbols are listed in Table 1.
(1) $Minimize\ \ \ \ T_{decode}$ $=T_{load}+T_{overlap}+T_{comp}$
(2) $M$ $\leq M_{max}$
The decode latency consists of three components: the first cross-layer-group loading time $T_{load}$, the cross-layer-group overlapping time $T_{overlap}$, and the final cross-layer-group computing time $T_{load}$, as in Eq. 1. The loading time $T_{load}$ is the weights missed in the cache divided by the flash loading bandwidth as $BW^{small}_{flash}$, as in Eq. 3. The final cross-layer-group computing time $T_{comp}$ is the group memory size $M_{cl}$ divided by the memory bandwidth $BW_{mem}$, as in Eq. 4. Furthermore, the overlapping time consists of two parts, i.e, the on-demand loading time $T_{load}$ and preloading latency $max(T_{preload},T_{comp})$, as in Eq. 5. We load the weights that are dissimilar across layers but not present in the cache, with latency $T_{load}$, as in Eq. 6. These weights typically have small chunk sizes, leading to lower bandwidth $BW^{small}_{flash}$. Preloading, on the other hand, loads weights at the cross-layer-group level, fetching only the cache-miss weights (Eq. 7). Since the chunk size in this stage is relatively large, the reading efficiency is significantly higher with bandwidth $BW^{large}_{flash}$.
(3) $T_{load}$ $=\frac{M_{cl}\cdot(1-hr)}{BW^{small}_{flash}}$
(4) $T_{comp}$ $=\frac{M_{cl}}{BW_{mem}}$
(5) $T_{overlap}$ $=T_{onload}+max(T_{preload},T_{comp})$
(6) $T_{onload}$ $=\frac{S_{l}\cdot(1-sp)\cdot(1-hr)\cdot(1-si)}{BW^{small}_{flash}}$
(7) $T_{preload}$ $=\frac{M_{cl}\cdot(1-hr)}{BW^{large}_{flash}}$
The memory cost also consists of three components: cross-layer group memory $M_{cl}$, weight cache memory $M_{cache}$, and KV cache memory $M_{kv}$ (Eq. 8). For the KV cache, we only consider the fixed-size case. Therefore, only the first two components will dynamically influence the memory cost. The cross-layer group memory is the size of active weights, as in Eq. 9.
(8) $M$ $=M_{cl}+M_{cache}+M_{kv}$
(9) $M_{cl}$ $=S_{l}\cdot(1-sp)\cdot N$
Preload-and-computation-balanced cross-layer group search. We determine the parameters ($sp$, $S_{cl}$, and $M_{cache}$) in a greedy manner, as follows. First, since LLM accuracy is only related to LLM sparsity, we set LLM sparsity by $sp=1-(M_{max}/S_{m})$ to ensure the highest accuracy. Second, minimize the decode time recursively. We increase layer number of cross-layer group $L$ in a step by step manner. This brings lower $T_{preload}$. In case $T_{preload}\leq T_{comp}$, then stop. Furthermore, if the $T_{preload}$ decrement is less than a threshold, then stop.
This approach ensures near-full memory utilization, minimal latency, and high accuracy. In case that the memory budget changes in online phase, we tune cache size to maintain well overlap between computation and flash read operations.
4.2. Dynamic LLM weight caching
To further reduce the number of loaded weights, we design the dynamic LLM weight caching based on observations of hot weights, as illustrated in Fig. 12. To maximize the cache hit rate, we track the frequency statistics of activation and evict the least-used weights in online phase.
To manage weight eviction, we maintain independent counters for the weights of each layer, ensuring a balanced cache size across all weights. If a newly activated channel has a higher count than the least-used channel in the cache, we evict the least-used channel. Fig. 12 illustrates an example of our dynamic cache mechanism. For a given sequence, we begin by initializing the usage count of all channels to zero. For the first token, channel index 0 is present in the cache, while channel indices 1, 4, and 6 need to be loaded from flash storage, resulting in a hit ratio of 25%. For the second token, channel indices 0, 4, and 6 hit in the cache, while only channel index 7 needs to be fetched from flash. Since channel index 1 has the lowest frequency, we replace it with channel index 7, improving the hit ratio to 75%.
Figure 12. An example of dynamic weights caching during LLM decoding. There are 8 channels in a weight but only half of channels are cached in memory.
5. Self-Distillation for Top-K Sparse LLM
Even with superior quality compared to other sparsity techniques, Top-K activation sparsity still introduce approximation for active weights selection. To compensate this, inspired by quantization-aware self-distillation, we propose Top-K sparsity-aware self-distillation. It has been integrated with the existing quantization self-distillation framework, together with the customer data (e.g., agent) fine-tuning as a standard step for actual LLM deployment on device, without adding extra overhead.
The reason to choose self-distillation is because traditional supervised fine tuning often fails to recover the model’s nuanced weight distributions and activation patterns, leading to performance loss. By comparison, self-distillation uses the full model’s output distribution as a soft target to preserve essential distributional details and weight correlations, thereby reducing errors and maintaining fine inference details for more robust, generalizable performance. It boosts accuracy with only a few to tens of GPU hours to learn on a few thousand samples, and generalizes across various sparsity levels.
Standard self-distillation faces two major challenges to apply for Top-K sparsity. First, gradient vanishing occurs because the sparsity mask introduces many zero values, causing most backpropagated gradients to also become zero, which slows down training. Second, loss mismatch arises because traditional cross-entropy loss does not effectively capture the relationship between the sparse LLM and the original model, leading to suboptimal distillation results.
Figure 13. The forward and backward process of self-distillation between Top-K activation sparsity and full weight model.
To overcome these issues, we propose top-k sparsity self-distillation as illustrated in Fig. 13, which introduces two key improvements:
The first is gradient straight-through estimation, a technique that reduces the impact of zero-masked values on gradient propagation, allowing more gradients to pass through and accelerating convergence. The second is the use of combine KLD(Kullback-Leibler Divergence)-CE(cross-entropy) loss, which better aligns the sparse LLM’s loss with that of the original model, improving the effectiveness of the distillation process.
5.1. Gradient STE and SD loss
Gradient Straight-Through Estimation. Sparsity leads to a large number of zero elements, which causes the gradients to be unable to update during backpropagation. To mitigate this, gradient straight-through estimation (STE) is employed. As shown in Equations 10 and 11, STE preserves the learning signal by replacing the gradient of the masking operation with an identity function during the backward pass:
(10) forward: $\quad y=\operatorname{Mask}(x)$
(11) backward: $\quad\frac{\partial y}{\partial x}=I$
By doing so, STE effectively bypasses the zero-masking effect in the gradient computation, allowing the gradients to flow as if the masking were transparent. This ensures that even in the presence of high sparsity, the model receives sufficient gradient information to update its parameters. Consequently, STE not only accelerates convergence by preventing gradient decay but also enhances the robustness of sparse inference training, thereby enabling the model to better retain and adapt the critical weight distributions and activation patterns.
Sparse-Distillation Loss. In our self-distillation framework, we denote $P_{T}$ as the output probability distribution of the teacher model and $P_{S}$ as that of the student model. We first define the Kullback-Leibler divergence as:
(12) $\mathcal{D}_{\text{KL}}(P\parallel Q)$ $=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$
In addition, we incorporate the cross-entropy loss between the teacher-provided labels $y_{T}$ and the student outputs $y_{S}$, denoted as $CE(y_{T},y_{S})$, which measures the discrepancy between the student prediction distribution and the teacher’s soft labels. Combining these two components, the self-distillation loss is formulated as:
(13) $\mathcal{L}_{\text{SD}}$ $=\gamma\,\mathcal{D}_{\text{KL}}(P_{T}\parallel P_{S})+(1-\gamma)\,CE(y_{T},y_{S}),$
where $\gamma\in[0,1]$ is a weighting factor that balances the two loss terms. It depends on sparsity ratio. Under high sparsity, $\gamma$ tends toward 0. Under low sparsity, it comes to 1. This occurs because, at high sparsity levels, sparse model’s output distribution deviates significantly from the original model. As a result, using KL divergence to measure the discrepancy becomes less reliable and may lead to performance degradation. In these cases, employing cross-entropy loss, which directly compares the predicted labels, provides a more effective training signal.
Minimizing $\mathcal{L}_{\text{SD}}$ encourages the student model not only to mimic the overall distribution of the teacher model but also to closely match its output values. This approach preserves crucial distributional information and weight correlations even under conditions of quantization and sparsity.
5.2. One-Distill-All-Scale: A Key Advantage of Distillation
A major advantage of our self-distillation approach is its capability of one-distill-all-scale. In conventional fine-tuning or distillation procedures, separate processes are often required for models at different sparsity levels, which can be both time-consuming and computationally expensive. In contrast, our method performs a single distillation at a high sparsity level, capturing the essential distributional characteristics and weight correlations of the original model.
Thanks to this process, the distilled student model can be effectively applied to scenarios with lower sparsity without additional fine-tuning. This is because the self-distillation not only transfers knowledge from the full teacher model but also establishes a robust framework that generalizes well across different sparsity scales. As a result, the model maintains consistent performance and efficiency, drastically reducing the overall training and deployment costs.
In summary, the one-distill-all-scale property enables our approach to streamline the distillation process and ensure adaptability across a range of sparsity levels, making it particularly attractive for efficient large-scale model inference.
6. Implementation
ActiveFlow is built on llama.cpp, a widely-used LLM inference framework for mobile devices. The whole model is stored in flash and only active weight, cached weight, and preloaded weight are in DRAM. This paper is based on the CPU backend of llama.cpp. The big cores execute computations and the little cores execute data loading concurrently. Since decoding speed is memory bandwidth bound, and mobile devices use a unified DRAM among all processors, we believe implementing ActiveFlow on different processors should have similar results. Past work (Wei et al., 2024; Wang et al., 2025b) have also demonstrated the superior performance of CPU over NPU for decoding on devices. We thus choose CPU in this paper for implementing convenience.
Flash loading. To implement cross-layer-group LLM weight loading, we modify the way weight tensors are stored in the GGUF format. Specifically, we save each operator’s weights as fundamental tensors organized in a cross-layer-group manner. We utilize IO uring, a low-overhead asynchronous I/O mechanism, to read the weights efficiently. In particular, we use the io_uring_prep_read and io_uring_submit functions to asynchronously request reads for active weights. After submitting all read requests, we synchronize the I/O operations using the io_uring_wait_cqe function. When reading active weights, we sparsely load different channels into a dense buffer, which helps optimize memory buffer layout for better compactness. Additionally, to ensure compatibility with quantization, we apply a transpose operation to the weights. This allows for complete retrieval of the necessary scaling factors when reading channels, thereby facilitating the quantization.
Swapping pipeline. To implement the active weight swapping pipeline, we first create a dedicated weight loading thread using the ggml_thread_create function. This thread is bound to a little core of the CPU via the sched_setaffinity function to optimize resource utilization. Synchronization between the weight loading thread and the main computing thread is achieved through atomic semaphores. We use atomic_load_explicit and atomic_store_explicit to manage a request signal and a complete signal that facilitates communication between the two threads. The signals operate at the cross-layer-group granularity, ensuring proper execution order between computing and weight loading operations.
Caching. Additionally, we implement the dynamic LLM weight caching, where caching is managed separately for each weight tensor. We use a hash table-based approach to efficiently query cached weight channels and dynamically track their activation frequency during decoding. When loading a new channel, we replace the least frequently activated channel, updating its index pointer in the hash table accordingly. Furthermore, we develop a kernel for generating active channel indices. This kernel maintains activation thresholds corresponding to different LLM sparsity levels. Before each activation step, it determines whether a channel should be activated based on the appropriate threshold.
Self-distillation. In order to implement and test sparsity-aware self-distillation, we extend BitDistiller (Du et al., 2024) into an open-source framework called Sparse-Distillation. We have developed a plug-and-play module that enables sparse model inference and training with just one line of code. This framework also supports quantization-aware sparsification. Specifically, we add a predictor to each block and perform activation sparsification by loading pre-generated block-level thresholds. During backpropagation, we incorporate a gradient straight-through estimation (STE) layer, and in the loss function design, we provide an SDloss option. In our self-distillation experiments, we train on a mixed dataset from alpaca-wikitext-c4, with each epoch containing approximately 9k data samples. A full training run comprises five epochs, with a learning rate of $8\times 10^{-6}$ and 4-bit quantization.
Overall, ActiveFlow comprises 3762 new lines of C++ code and thousands lines of Python code.
7. Evaluation
Table 2. The hardware devices for evaluation.
Device CPU Memory Flash (MaxBW)
OnePlus 12 X4+A720+A520 16GB UFS 4.0 (5.8 GB/s)
Pixel 6 X1+A76+A55 8GB UFS 3.1 (4.2 GB/s)
Infinix ZERO 30 A76+A55 8GB UFS 2.2 (3.6 GB/s)
Figure 14. The end-to-end decoding speed, perplexity and memory cost of three LLMs compared with baselines on various devices. Each point represents a sparsity ratio: from left to right 0.8, 0.7, 0.6, 0.5. Since decoding is memory bound, latency increases with less sparsity and more memory cost.
We evaluate ActiveFlow on both end-to-end and technique performance, compared to several baselines. The evaluation setup is as follows:
7.1. Evaluation setup
Hardware devices. As shown in Table 2, we evaluate ActiveFlow on three mobile devices, covering a range from high-end to low-end. For clarity, we label the three devices as Device 1, Device 2, and Device 3.
Models. To assess end-to-end performance, we test popular LLMs, including the Llama and Mixtral series, with model sizes ranging from 7B to 56B parameters. All LLMs undergo 4-bit quantization using Q4_0, a widely used technique that has minimal impact on accuracy. For the technique evaluation, we extract and use eight layers from the original LLM.
Baselines. We compare ActiveFlow against llama.cpp in terms of decoding speed and memory usage. For perplexity and accuracy evaluation, we use the original LLM, ProSparse, and TEAL as baselines. ProSparse and TEAL represent state-of-the-art ReLU-sparse and top-k-sparse LLMs, respectively.
Measurement. Our evaluation focuses on decoding speed, perplexity, accuracy, latency, hit rate, memory cost, power, and energy consumption. We use the clock_gettime function to record start and end timestamps, computing latency as the difference between them. We measure the total number of decoded tokens and the total decoding time, calculating speed as $N_{tokens}/Latency$. We use lm-eval-harness, a widely used LLM evaluation framework, to measure perplexity on the WikiText-2 tasks. We track cache hits and misses, computing the hit rate as $N_{hit}/(N_{hit}+N_{miss})$. We analyze memory cost using the Android Studio Profiler. We obtaine current and voltage values by reading system files (voltage_now and current_now) to calculate power consumption. These values are collected every 0.5 seconds on average, and we use the decoding latency to compute the overall energy consumption.
7.2. End-to-end performance
Decoding speed. We first evaluate the decoding speed of different LLMs across various devices under different memory cost conditions as illustrated in Fig. 14(a). For Device 2 and Device 3, using the LLaMA-2-7B model, we achieved the same performance as the full-weight memory setting while reducing memory cost by 40%. When reducing memory cost by 75%, our method achieved a 1.9× and 1.5× speedup compared to the full-weight in-memory setting on Device 2 and Device 3, respectively. The speedup is primarily due to our computing-loading pipeline, which enables higher decoding speed even under lower memory cost constraints. However, on Device 1, when using 60% of the memory cost, our performance dropped by 54% compared to the full-weight memory setting. This is because the CPU compute bandwidth of Device 1 is significantly higher than its flash read bandwidth, making the pipeline constrained by flash bandwidth. Nonetheless, at 75% memory cost, our method was able to achieve a decoding speed of 5.9 tokens per second.
For the Mixtral model, we successfully enable decoding under 6GB of memory. When the memory cost was 4.3GB, the decoding speed on Device 1, Device 2, and Device 3 was 1.3, 1.0, and 0.4 tokens per second, respectively. As the memory cost was reduced to 2.9GB, the performance improved to 2.3, 1.8, and 0.8 tokens per second, achieving a 1.8× to 2.0× speedup across the three devices.
Perplexity. We evaluate the perplexity and accuracy of different LLMs under varying memory cost conditions as illustrated in Fig. 14(b). For the LLaMA-2-7B model and LLaMA-3-8B model, our method achieved results comparable to the full-weight memory setting at 60% memory cost. For Mixtral-8x7B, we achieve a perplexity comparable to the baseline (24.6GB) while using only 4.4GB of memory. With further memory reduction, perplexity increases due to the increased sparsity of the LLM, which leads to a decline in inference capability.
7.3. Technique breakdown
To validate the effectiveness of our system’s techniques, we conduct ablation studies and standalone tests for each component, evaluating their impact on decoding speed, perplexity, and hit rate.
Cross-layer-group pipeline. First, we examine the effect of the cross-layer-group pipeline on decoding speed, as shown in Fig. 15. We used a 60% sparsity LLaMA-2-7B model and tested it across three devices. Our baseline consisted of serial computation and memory reads. Experimental results show that when the layer number in a cross-layer group is set to 1, the average speedup across all three devices is 10%. However, increasing the layer number to 4 results in a 120% performance improvement, as it enhances the efficiency of flash memory reads. Finally, with the addition of Dynamic Cache, our method achieves 2$\times$, 2.3$\times$, and 3$\times$ speedups over the baseline on the three devices, respectively.
To further understand the benefits and overhead of each technique, we conducted individual experiments for detailed analysis. As shown in Fig. 16, we evaluated the trade-offs of cross-layer loading. In Fig. 16(a), we measured the loading and preloading overhead for a single layer when the layer number in a cross-layer group is set to 1, under different cosine similarity values. The results show that when cosine similarity is lower than 0.2, the preload latency is lower than the on-demand load latency. However, when cosine similarity exceeds 0.4, the on-demand load latency becomes lower than the preload latency. Since the cosine similarity of most layers is above 0.8, our cross-layer approach effectively overlaps preloading and computation, optimizing performance.
Figure 15. The decode speed improvement of LLaMA-2-7B model on three devices by each technique.
Figure 16. The performance and memory cost of cross-layer loading.
In Fig. 16(b), we evaluate an 8-layer decoder of LLaMA-2-7B, measuring preload, load, and total latency as well as memory cost under different layer numbers in a cross-layer group. When the layer number is 0, computation and flash loading occur sequentially, leading to high total latency. When the layer number increases to 1, computation begins to overlap with preloading, reducing total latency by 52%. As the layer number further increases to 4, improved preload efficiency enables a 4.1× speedup compared to the size 0 setting. However, increasing the layer number also leads to higher memory cost, introducing additional overhead. Overall, increasing the layer number in a cross-layer group effectively enhances decoding performance, while the additional memory overhead remains relatively low.
Contextual caching policy. In Fig. 17, we compare the performance of context-level cache and task-level cache, where the task-level cache is built based on hot weight statistics from the WikiText-2 dataset. Fig. 17(a) presents the cache hit rate of context-level and task-level caches under different token lengths, evaluated on a sequence of BoolQ dataset. When the token length is 10, the context-level cache achieves a 13% higher hit rate than the task-level cache. As the token length increases to 40, the context-level cache’s hit rate drops from 77% to 74%, yet it remains 10% higher than the task-level cache. Fig. 17(b) shows the cache hit rate across different downstream tasks. The task-level cache’s hit rate varies between 54% and 74% across different datasets due to variations in hot weight distributions among tasks. In contrast, our context-level cache effectively adapts to these variations, achieving an average hit rate improvement of 12% over the task-level cache.
Figure 17. The performance of task-level and context-level cache.
Figure 18. The perplexity of baseline and top-k sparsity-aware self-distillation.
Figure 19. The power and energy consumption of ActiveFlow and baseline.
Top-k sparsity-aware self-distillation. Fig. 18 shows the perplexity changes of baseline and our top-k sparsity-aware self-distillation. Compared with the baseline, our approach substantially lowers the perplexity, especially when the sparsity exceeds 80%, resulting in remarkable improvements in LLM accuracy. The improvement in accuracy can be attributed to the incorporation of the gradient straight-through estimator and the adaptive combined loss function in our self-distillation process.
7.4. Power and energy consumption
We also evaluate the power consumption and energy efficiency of ActiveFlow compared to the baseline on Device 1, as shown in Fig. 19. During the decoding process, ActiveFlow’s average power consumption is 1.86W, which is 27.34% lower than llama.cpp. This reduction is primarily due to computation wait time while fetching data from flash memory in our overlap pipeline. However, since ActiveFlow reduces the required weight to compute and load, the processing time per token is also reduced. As memory cost decreases, the energy consumption per token further declines. Specifically, at a 1.3GB memory cost, our system achieves a 53% energy reduction compared to llama.cpp.
8. Related Works
Sparsity in LLMs. Sparsity in LLMs have been the focus of many research efforts. Mirzedeh et al. (Mirzadeh et al., 2023) propose replacing the ReLU activation function in LLMs to reduce computation and weight transfer. HiRE (L et al., 2024) introduces high-recall approximate Top-K estimation for sparse LLMs. Prosparse (Song et al., 2024a) leverages the sparsity of the ReLU function and the gated branches within the feed-forward network (FFN) to predict model sparsity. Q-Sparse (Wang et al., 2024c) is an effective approach to training sparsely activated LLMs from scratch. TEAL (Liu et al., 2025a) is a training-free method that applies magnitude-based activation sparsity to LLM activations. Compared to these works, ActiveFlow further optimizes LLM weights through self-distillation, taking into account quantization and sparsity, and achieves SOTA scaling law on accuracy in both generation and downstream tasks.
Efficient LLM inference system. Based on the sparse activation characteristics of LLMs, previous work has designed systems for efficient model loading and inference. DeJavu (Liu et al., 2023) is a system that employs a low-cost algorithm to dynamically predict contextual sparsity. Alizadeh et al. (Alizadeh et al., 2024) propose a system that enables effective inference of LLMs on devices with limited memory. PowerInfer (Song et al., 2024b) is a CPU-GPU hybrid inference engine. ActiveFlow stands out from these works by (1) supporting any LLMs without the limitation of the ReLU activation function, and (2) providing elastic memory management for various devices with limited memory.
Static pruning techniques. Static pruning and quantization methods have been widely explored for compressing large language models. For instance, CFSP (Wang et al., 2025a) leverages structural redundancy for efficient pruning, while DB-LLM (Chen et al., 2024) and PB-LLM (Shang et al., 2023) propose dynamic block pruning strategies to reduce model parameters. In addition, RIA (Biswas et al., 2025) introduces regularization-based approaches to enhance quantization accuracy. These techniques enable substantial reductions in model size and computation without extensive retraining. However, they necessitate additional processing of model weights, making dynamic task adaptation and flexible loading challenging. ActiveFlow is nor only compatible with static weight optimize techniques but also supports dynamic processing.
Knowledge distillation for LLMs. Quantization-aware distillation has emerged as an effective paradigm to compress LLMs while mitigating accuracy loss. BitDistiller (Du et al., 2024) exemplifies this by integrating low-bit quantization with teacher-student distillation, achieving competitive performance under tight resource constraints. Other works (Zhao and Zhao, 2024) further refine the alignment between teacher and student models under quantization, balancing compression and task performance. Such approaches complement dynamic pruning methods and are especially attractive for deployment on devices with limited memory. ActiveFlow considers both quantization and sparsity factors during self-distillation, achieving SOTA in accuracy for both generation and downstream tasks.
9. Conclusion
This paper proposes the first LLM inference system on mobile devices that supports adaptive DRAM usage, in order to scale up the deployable model size. It is based on the idea of active weight swapping between DRAM and flash, integrating three novel techniques: cross-layer active weight preloading, sparsity-aware self-distillation, and active weight swapping pipeline. It achieves the inference performance-cost Pareto frontier compared to other efficiency optimization methods. This paper breaks the DRAM limitation for LLM deployment, opening up the new opportunity of server-level LLMs deployment on mobile devices.